5. The reading given by a thermometer in ice water (actual temperature 0° C) is a random variable with pdf otherwise where k is a constant. (a) Find k b) Find the probability that the reading is within 25 (e) Find the expected reading. (d) Find the median reading. (e) Find the standard deviation of the reading. C of the actual temperature
Answer #1

a) for this to be valid: int_{-1}^{2} f(x)dx must be 1.

int_{-1}^{2} f(x)dx =int_{-1}^{2} k*(4-x2) dx =k(4x-x3/3) |2-1 =k*9=1

k=1/9

b)

P(-0.25<X<0.25)=с 0.25 f(x)dx =с 0.25 (1/9)*(4-x2) dx =(1/9)(4x-x3/3) |0.25-0.25 =0.221065

c)

expected value =int_{-1}^{2} xf(x)dx =int_{-1}^{2} (1/9)*(4x-x3) dx =(1/9)(4x2/2-x4/4) |2-1 =0.25

d)

let median be at x=m

hence int_{-1}^{m} f(x)dx =int_{-1}^{m} (1/9*(4-x2) dx =(1/9)(4x-x3/3) |m-1 =(1/9)*(4m-m3/3)-(-0.40741)=0.5

solving above median=0.2091

e)

E(X2)=int_{-1}^{2} x2f(x)dx =int_{-1}^{2} (1/9)*(4x2-x4) dx =(1/9)(4x3/3-x5/5) |2-1 =0.6

hence std deviation =sqrt(E(X2)-(E(X))2) =sqrt(0.6-0.252) =0.7331